Hello,
I'm seeking help with the following math problem below, any help would be appreciated.
(1/2)a + (3/4)b - c
when a=8, b=4 and c=-2
Thank You!

Answer by ewatrrr(24785)


Hi,
(1/2)a + (3/4)b - c when a= 8, b= 4 and c= -2
(1/2)*8 +(3/4)*4 -(-2)
4 + 3 + 2 = 9